OpenStudy (denise):

5(4/5)^2+8(4/5)-6 =???

OpenStudy (anonymous):

18/5

OpenStudy (denise):

how did you get that

OpenStudy (anonymous):

see 5 (4/5)^2 is 5 (16/25) = 16/5

OpenStudy (anonymous):

8(4/5) = 32/5

OpenStudy (anonymous):

u add both f these u get 48/5

OpenStudy (anonymous):

then u subtract 6 which is (48/5) - 6 (48-30)/5 18/5

OpenStudy (anonymous):

got it?

OpenStudy (denise):

thanks! I am trying to do relative minima and maxima problems

OpenStudy (denise):

I am lost as hell!

OpenStudy (anonymous):

how old r u?

OpenStudy (denise):

40

OpenStudy (denise):

just went back to college

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

ur my mums age almost

OpenStudy (denise):

haha

OpenStudy (anonymous):

cn i help smwher?

OpenStudy (denise):

that would be awesome

OpenStudy (denise):

how do you determine when it is a minima or a maxima

OpenStudy (anonymous):

see lets take an example of a cubic function f(x)

OpenStudy (anonymous):

when you differentiate it to get f'(x) , u evaluate the roots of f'(x)

OpenStudy (anonymous):

these r called critical points...these are the points where it can be maximum or minimum..the function that is

OpenStudy (anonymous):

get it til here?

OpenStudy (denise):

k.

OpenStudy (denise):

got that

OpenStudy (anonymous):

so now to determine whether its max or min, what u do is u differentiate once more to get f''(x)..now plug in ur critical point into it

OpenStudy (denise):

Are you going to be on here long today? I have about 60 problems I have to do.

OpenStudy (denise):

What determines if it is maxima or minima?

OpenStudy (anonymous):

when u plug the point into f''(x), see the sign of the function

OpenStudy (denise):

if its negative its min and positive max?

OpenStudy (anonymous):

if the sign of f"(x) is negative its maxima, and viceversa

OpenStudy (anonymous):

no the opposite

OpenStudy (denise):

gotcha! Thanks a million!

OpenStudy (anonymous):

m not gonna b long inuf fr 60 problems..bt as much as i can

OpenStudy (denise):

thanks I am afraid I will be on here all weekend

OpenStudy (anonymous):

no prob...well see wt we cn crack

OpenStudy (denise):

According to th answer on mymathlab, the answer was -46/5.... how is that

OpenStudy (amistre64):

nuthin wrong with 40s :)

OpenStudy (anonymous):

well amistre check on that question she posted

OpenStudy (amistre64):

which one?

OpenStudy (anonymous):

right on top..

OpenStudy (denise):

I really appreciate yall helping me!!!1

OpenStudy (amistre64):

5(4/5)^2+8(4/5)-6 =??? this one

OpenStudy (denise):

yep

OpenStudy (anonymous):

yeah

OpenStudy (amistre64):

5x^2 +8x -6 is a parabola; that looks like a bowl sitting on the counter; it has a local minimum at its vertex; ....i sctually finished typing this reply like 40 seconds ago; slow response time i guess

OpenStudy (amistre64):

-b/2a = x compnent; -8/2(5) = -4/5 right?

OpenStudy (amistre64):

5(-4/5)^2+8(-4/5)-6 = y ^^^ use (-)4/5

OpenStudy (amistre64):

5(16/25) +8(-4/5)-6 = y

OpenStudy (anonymous):

yeah denise i thought it was 4/5 while actually its -4/5

OpenStudy (amistre64):

16 -32 -6 --------- = 5 5 1

OpenStudy (amistre64):

6(5) = 30 s000 (16-32-30)/5

OpenStudy (amistre64):

-46/5

OpenStudy (amistre64):

makes sense?

OpenStudy (anonymous):

yeah

OpenStudy (denise):

k

OpenStudy (denise):

where did the 6(5) come from???

OpenStudy (anonymous):

cz u take the lcm while subtracting

OpenStudy (denise):

to get the 30

OpenStudy (amistre64):

yeah; fractions reguire you to be wearing the same pants; and that 6/1 is not dressup for company; so we got to changes its clothes to match .... we put it in 5 pants and find a top to match

OpenStudy (denise):

lol!